# Demystifying Behavior-Based Malware Detection at Endpoints

### v1

@@ 7.4. Combatting Environment-Sensitivity With Environment-Invariant Features
We have shown that the environment-sensitivity of program behaviors impacts the model. When the classifier exploit features that are specific to the training environment (e.g., trace length), its predictions fail to generalize. Table 10 suggests that the artifacts of SB1 are absent in SB2, and vice versa, as they are configured sufficiently differently. This highlights an opportunity: learning features that are invariant across two environments (e.g., both sandboxes) might yield better generalization to endpoint detection.
@@
Prior work has studied invariant learning in self-supervised learning, e.g., viewpoint-invariant visual features (purushwalkam2020demystifying, ). We employ the Siamese loss (chen2021exploring, ), which forces the model to produce similar embeddings (measured by a distance metric) on pairs of semantically related inputs. In our case, each pair consists of two traces of the same sample from different sandboxes (in SB->EP) or from different EP hosts (in EP->EP). To gain robustness to variable behaviors, we aim to make the model invariant to the differences in these pairs (such as sandbox-specific artifacts). The critical question is whether invariance across SB traces translates to invariance across EP traces. We represent a pair of traces of a sample $P_{i}$ as ($x_{i,0},x_{i,1})$ and the set of all pairs as $\mathbf{D}$; and define the following loss function that is added to the standard ERM loss as a regularizer:
@@
Table 12. The impact of $\mathcal{L}_{\text{inv}}$ on endpoint detection TPR.
@@
Figure 5. The model’s embeddings with increasing values of $\alpha$ for invariant learning. Sep quantifies the separability of the SB1 and SB2 embeddings. t-SNE (van2008visualizing, ) visualization.
@@
8. Endpoint Malware Detection Leaderboard
Our investigation has relied on a dataset of endpoint traces from the wild provided to us by Avllazagaj et al., who collected and analyzed it in their prior work (avllazagaj2021malware, ). We have pinpointed the salient challenges in ML-based endpoint malware detection using this data. Performing evaluations only on lab-based data can obscure these challenges and produce biased solutions inapplicable to the real world (cherubin2022online, ; jacobs2022ai, ). To minimize this bias, we designed a leaderboard for realistic evaluations of behavior-based malware detectors (malwaredetectioninthewild.github.io (leaderboard, )). Prior leaderboards such as RobustBench (croce2021robustbench, ) or TrojAI (trojai, ) have motivated research into critical and difficult problems and fueled many innovations in ML.
Anyone may submit a pre-trained detector and a feature extractor that converts a trace in the standardized format into an input to the detector. We evaluate each submitted detector on our endpoint traces and list it on our leaderboard. Although we are unable to share Avllazagaj et al.’s dataset, we will release the sandbox dataset we collected and the metadata for over 200K samples used in our work to assist the participants. The details of these artifacts can be found in §E. This format offers a blueprint for security vendors to guide researchers without publicly sharing their sensitive data. We hope to spearhead it into more applications of ML for security.
9. Discussion
@@
Threats to Validity and Limitations. Our study has some limitations due to the nature of our endpoint data. First, this data was collected at the hosts of a single security vendor. Although we cannot rule out selection bias, the fact that these hosts are located in over 100 countries in both enterprise and consumer settings suggests that our results have broad applicability. Second, our data was collected five years ago. This is a common limitation in malware detection studies as collecting large-scale up-to-date data has not been feasible for researchers (yang2021cade, ; barbero2022transcending, ; chen2023continuous, ). We focus on the gaps between sandbox-based and endpoint detection and not on the specifics of the threats that existed at the time, making our observations relevant to today’s malware detection landscape. Third, our data consists of only Windows hosts and lacks any network-related actions. As 95% of malware is aimed at Windows (windowsmalware, ) and ML-based detectors can perform well without network actions (jindal2019neurlux, ), we do not expect this to affect our findings.
@@
Early Detection of Malware. In our study, a detector uses the whole execution trace of a sample that terminates in a host. Although vendors offer ways to undo the damage from malware after its execution, e.g., quarantining (quarantine, ), it is more desirable to catch malware in its tracks. This would allow for containing the harm of a malware sample by cutting its execution short, e.g., before a ransomware sample starts encrypting personal files. As the detector has strictly less information available in this task, our results can be viewed as an upper bound on the performance of early malware detection. In consequence, we do not expect a detector that is trained only on sandbox traces to perform well in this setting.
@@
Fine-Tuning a Sandbox-Based Model on Endpoint Traces. In Appendix D, we present an experiment where we first train a NGR model on sandbox traces and then fine-tune this model on increasing amounts of data from our EP training set. This hybrid approach simulates a realistic scenario where low-cost sandbox data is supplemented with high-cost endpoint traces from the wild. We evaluate two strategies: fine-tuning all layers of the model and fine-tuning only the last layer. We find that adding a few EP traces improves performance, and when very few EP traces are available, fine-tuning outperforms training from scratch (3–5% higher TPR).
@@
Attacks on ML. All ML-based detectors, especially deep learning ones, are subject to adversarial attacks due to their sensitivity to input perturbations (szegedy2014intriguing, ; grosse2017adversarial, ). We expect our models will also be vulnerable to such attacks, e.g., an adversary can inject dummy actions into the behaviors of their malware sample designed to fool a model. We consider a defense against these attacks to be out of scope for this work; instead, we aim to show that naturally occurring factors, such as the environment variability, also degrade the performance of ML-based detectors in the wild.
10. Conclusion
@@
Malware detection with dynamic features serves as a last line of defense at endpoints, providing security when all other measures have failed. Malware samples that breach this last line cause real-world infections and harm. A long-standing ambition is to find the best way to detect such samples. Through a systematic exploration of different scenarios, we provide some clarity. Existing detectors trained on sandbox executions degrade when deployed at endpoints. Training on real-world behaviors performs much better, though it necessitates collecting data from real hosts. We pinpoint various challenges in endpoint detection and propose ML techniques targeting them, yielding moderate improvements. Ultimately, this task remains challenging, and through our leaderboard, we hope to guide the community to develop effective and applicable solutions.
11. Acknowledgements
@@
This research was supported by the Intelligence Community Postdoctoral Fellowship, the U.S. Department of Defense, EPSRC Grant EP/X015971/1, National Science Foundation Grant CNS-2154873, and generous gifts from Google and Amazon. We would like to thank VirusTotal for granting us academic access to their platform, which allowed us to curate and label our datasets. Finally, we thank Omer Faruk Akgul and David Acs for their support in the earlier phases of our research and Erin Avllazagaj for providing us with access to their program traces in the wild dataset. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the supporting organizations.
@@ Appendix A Trace Standardization
First, we only keep the action types that all formats in our datasets share, which are file creation, registry key creation and deletion, process creation and injection, and mutex creation. Second, we clean up the strings (e.g., file names) in each trace by removing white spaces, capitalization, punctuation, non-ASCII characters, and so on. Third, we replace certain file and directory names to minimize differences caused by operating system versions or logging conventions. Fourth, we use regular expressions to replace MAC addresses, Windows security and resource identifiers, hashes (SHA-256, SHA-1, MD5), and epoch timestamps with special tokens (e.g. <macaddress>), to prevent introducing artifacts to our ML models. Fifth, we tokenize the entries in each trace (e.g., split a full file path into its components) and save a trace as a sequence of tokens, following (jindal2019neurlux, ). This standardized format allows us to implement a leaderboard for endpoint malware detection (detailed in Section 8) by providing all participants with an expected input format that their detectors should accept. When we train a model, we only keep the top-10K tokens in terms of frequency in the training traces and replace the remaining tokens with special tokens, such as <rare_file_name>, again following (jindal2019neurlux, ; trizna2023nebula, ). This makes the task more suitable for learning by eliminating uninformative tokens and reducing the feature dimensionality.
@@ Appendix B Further Machine Learning Details
NGR (based on Maldy (karbab2019maldy, )). We convert each trace into a list of 2-grams. For example, the file path a/b/c.jpg is turned into three 2-grams: <a/b>, <b/c> and <c.jpg>. As this results in an intractable number of unique 2-grams (mostly very rare), we apply the hashing trick (weinberger2009feature, ) that assigns a numerical value up to $2^{14}$ to each 2-gram. The final feature vector—$x$—for a trace is a $2^{14}$-dimensional vector and each dimension is set to the number of occurrences of the corresponding 2-gram in the trace. We use a ResNet-based architecture (gorishniy2021revisiting, ) to train on these features.
HYB (based on Neurlux (jindal2019neurlux, )). We treat a trace as a natural language document and train an attention-convolution-hybrid sequence classification model. This approach eliminates the need for feature engineering (unlike the n-gram approach) and can extract useful features from long sequences thanks to the attention mechanism.
ATT (based on Nebula (trizna2023nebula, )). We treat a trace the same way as HYB (a sequence) and train a self-attention transformer-based architecture that is claimed to be robust to heterogeneous information (  e.g., different report format).
@@
Appendix C Case Studies on Sandbox-Specific Artifacts
@@
Here, we dive deeper into the sandbox-specific artifacts we identified in Table 10.
@@
SogouExplorer is a Chinese web browser that exists only in malware traces (100% MalR.) in SB1; whereas it does not exist in any SB2 traces and very few EP traces. The samples that interact with it mostly belong to families such as Sivis and Memery, all tagged as file infectors that attach their code to other programs. Considering that SB1 was developed by a Chinese vendor, we believe that they pre-installed this browser on their sandboxes to generate an analysis environment representative of Chinese hosts. This, however, creates features specific to SB1 as samples interact with the programs in the environment. Although this artifact exists in a few endpoint traces from hosts in China, its prevalence is almost zero.
@@
PersonalBankPortal, according to our research, is a program distributed by a Chinese bank to its customers. The samples that inject into this program belong to families such as Tinba and Ramnit, all considered as banking trojans that specifically ex-filtrate banking data. We believe the vendor pre-installs this program to lure malware samples into exhibiting their behaviors. Although this practice is useful for analyzing a sample (yong2021inside, ) (and for SB->SB), it causes artifacts that are rarely observed in the wild.
@@
Among the artifacts found in SB2, Spotify is a popular music streaming service, and Python is the interpreter for Python programming language. Both programs are targeted and injected by file infectors, similar to SogouExplorer in SB1. These programs are much less prevalent in endpoint traces than in SB2 traces. We believe the SB2 vendor, which is based in the US, pre-installs them to create an environment representative of the hosts in the US.

### v2

@@ VII-D Invariant Learning to Counter Environment-Sensitivity
We have shown that the environment-sensitivity of program behaviors impacts the model. When the classifier exploits features that are specific to the training environment (e.g., a sandbox artifact), its predictions fail to generalize. Table IX suggests that the artifacts of SB1 are absent in SB2, and vice versa, as they are configured sufficiently differently. This highlights an opportunity: learning features that are invariant across two environments (e.g., both sandboxes) might yield better generalization to endpoint detection.
@@
Prior work has studied invariant learning in self-supervised learning, e.g., viewpoint-invariant visual features [79]. We employ the Siamese loss [80], which forces the model to produce similar embeddings (measured by a distance metric) on pairs of related inputs. In our case, each pair consists of two traces of the same sample from different sandboxes (in SB->EP) or from different EP hosts (in EP->EP). We aim to make the model invariant to the differences in these pairs (such as sandbox-specific artifacts) to gain robustness to variable behaviors. The critical question is whether invariance across SB traces translates to invariance across EP traces. We represent a pair of traces of a sample $P_{i}$ as ($x_{i,0},x_{i,1})$ and the set of all pairs as $\mathbf{D}$; and define the following loss function that is added to the standard ERM loss as a regularizer:
@@
TABLE X: The impact of $\mathcal{L}_{\text{inv}}$ on endpoint detection TPR.
@@
Figure 4: The model’s embeddings with increasing values of $\alpha$ for invariant learning. Sep quantifies the separability of the SB1 and SB2 embeddings. t-SNE [81] visualization.
@@
Takeaways: An execution environment’s specific configuration might introduce artifacts into program traces, hurting behavioral detectors’ generalizability to other environments. Learning invariant features over a sample’s traces from multiple environments (e.g., two sandboxes) can enhance ML models’ robustness to such artifacts, improving their performance in real-world hosts with diverse configurations.
VIII Discussion and Limitations
@@
Dataset Bias, Size and Label Noise. Our study has some limitations due to our endpoint data. First, this data was collected from hosts that use a single vendor’s anti-malware product. Although we cannot rule out selection bias, the fact that these hosts are located in over 100 countries in both enterprise and consumer settings suggests that our results have broad applicability. Second, our data was collected six years ago. This is a common limitation in malware studies as collecting large-scale, up-to-date real-world data is infeasible for researchers [60, 25, 39]. Nevertheless, to our knowledge, our endpoint dataset (originally collected in [15]) is the only dataset of endpoint malware behavior analyzed in the literature. We focus on the gaps between sandbox-based and endpoint detection and not on the specifics of the threats of the time, making our observations still relevant today. Third, our data consists of only Windows hosts and lacks network-related actions. As 95% of malware is aimed at Windows [82] and ML-based detectors can perform well without network actions [6], we do not expect this to affect our findings.
Moreover, our endpoint dataset is much smaller than our sandbox dataset and smaller than what is available to a security vendor, potentially causing us to underestimate the EP->EP performance in practice. However, consider Figure 6 where we train models in EP->EP using increasing proportions of our endpoint dataset. Here, the performance resulting from using 25%, 50%, and 100% of the data for training is 35%, 41%, and 46% TPR, respectively. This hints that even with larger endpoint dataset sizes, the performance in EP->EP is unlikely to catch SB->SB due to fundamental challenges we exposed (difficult-to-classify samples and behavior variability).
Finally, for a given sample, labels obtained from VirusTotal (VT) are initially noisy and stabilize after around a year [83]. In our experiments (see Section IV-A), we use the oldest label available (e.g., from old VT reports) for the training samples and the labels from VT reports collected four years later (in 2022) for the testing samples. This methodology approximates the realistic conditions for deploying a malware model. To gauge the impact of label noise on our measurements, we compare the labels of training samples from recent VT reports to the labels we used to train our models. In the SB and EP training sets, old and new labels agree on 98% and 97% of the samples, suggesting that the effect of label noise is minimal.
@@
Early Detection of Malware. We study detectors that use the whole execution trace of a sample, and thus can only detect malware once it terminates. Although vendors offer ways to undo the damage from malware after its execution, such as quarantining [8], it is more desirable to catch malware in its tracks, e.g., before a ransomware sample starts encrypting personal files. In that task, a detector has strictly less information available to classify a sample, which makes our results an upper bound on the performance of early malware detection.
@@
Fine-Tuning a Sandbox-Based Model on Endpoint Traces. In Appendix F, we present an experiment where we first train a NGR model only on sandbox traces; we then fine-tune this model on increasing amounts of data from our EP training set. This hybrid approach simulates a realistic scenario where low-cost sandbox data is supplemented with high-cost endpoint traces from the wild. We implement two strategies: fine-tuning all layers of the model and fine-tuning only the last layer. We find that adding a few EP traces improves performance, and when very few EP traces are available, fine-tuning outperforms training from scratch (3–5% higher TPR).
@@
Attacks Against ML. All ML-based detectors, especially deep-learning-based ones, are subject to adversarial attacks due to their sensitivity to input perturbations [84, 85]. We expect our models will also be vulnerable to such attacks, e.g., an adversary can inject dummy actions into the behaviors of their malware sample designed to fool a model. We consider a defense against these attacks to be out of scope for this work; instead, we aim to show that naturally occurring pressures, such as environment variability, also greatly degrade the performance of ML-based malware detectors in the wild, posing a critical security and trustworthiness challenge in practice.
Promoting Future Research. Our investigation relies on a dataset of endpoint traces from the wild provided by Avllazagaj et al., who collected and analyzed it in their prior work [15]. Using this data, we have pinpointed the salient challenges in ML-based endpoint malware detection. Performing evaluations only on lab-based data can obscure these challenges and produce biased or unreliable solutions inapplicable to the real world [44, 43]. Although its terms of use prevent us from sharing Avllazagaj et al.’s dataset, we built a pipeline that allows the community to perform realistic evaluations of ML-based behavioral malware detectors, aiming to minimize this bias. Anyone may submit to us a pre-trained detector and a feature extractor that converts a trace in the standardized format (Section IV-A) into an input to the detector. Upon receiving a submission, we will evaluate it on our endpoint traces dataset and return the submitter a detailed performance report to guide their research. Further, if the submitter wishes, we will publicly list their results and submission details on the leaderboard on our website (https://malwaredetectioninthewild.github.io). To assist participants, we will release the sandbox dataset we collected and the metadata for over 200K samples used in our work. The details of these artifacts can be found in Appendix G. Please refer to our website for technical details, data access requests, and detector submission instructions. This pipeline offers a blueprint for security vendors to drive research without publicly sharing their sensitive data. We hope to spearhead it into more applications of ML for security where evaluation problems frequently hinder real progress.
Implications for Future Work. As described in Section III-A, malware detection in practice involves a pipeline of techniques. We believe ignoring this pipeline and treating each component as a standalone solution causes critical evaluation problems. Consider ML-based static malware detectors, an active research area [49, 4, 26]. They are ideally evaluated after filtering out the test samples that can already be detected by the pipeline components preceding static ML, such as existing blocklists, allowlists, and static signatures. However, to our knowledge, this is not an established practice, which might cause an overestimation of real-world performance (similar to our findings in Section VI). Our endpoint data allowed us to isolate and realistically evaluate ML-based behavioral detectors, revealing previously unknown challenges. This, however, still falls short of an ideal end-to-end evaluation. Instead of measuring the success of individual components, quantifying the overall effectiveness of a pipeline yields more relevant insights for users. For example, if static components detect 95% of the active malware at a given time and the dynamic components detect 80% of the remaining malware, the overall pipeline would be at 99% TPR. Moreover, the inter-dependence in components introduces poorly understood challenges, e.g., the trade-off between minimizing the pipeline’s false positives and classifying more samples statically to reduce infections. Due to our dataset limitations, we are unable to perform such analyses. We recommend that future research approach malware detection holistically, treating it as a pipeline, and collect suitable datasets for performing end-to-end evaluations.
Furthermore, prior work on ML-based behavioral detectors has focused strictly on a sandbox-based scenario where training and testing traces are collected from the same sandbox. This obscures the impact of sandbox configurations on models’ robustness when deployed on traces from diverse environments, such as real-world hosts, as shown in Section VII. Our work had to rely on traces from two third-party sandboxes with unknown configurations, which prevented us from conducting controlled experiments to study this impact, e.g., to increase the performance in SB->EP. Instead, we show that invariant learning over traces on these two sandboxes might increase robustness. While sandbox configuration is a well-studied topic in malware analysis [35, 21, 22], its role in generating training sets for ML-based detectors has been overlooked. To address this gap, we recommend that future work tackle realistic scenarios where behavioral models are deployed in unpredictable environments, emphasizing the need for robustness. Developing principled methods for configuring sandboxes for this scenario is a promising research direction.
IX Conclusion
@@
Behavioral malware detection serves as a last line of defense at endpoint hosts, providing security when all other measures have failed. Malware samples that breach this last line cause real-world infections and harm. A long-standing ambition is to find the best way to detect such samples. Through a systematic exploration of different scenarios, we provide some clarity. State-of-the-art ML-based detectors trained on sandbox execution traces degrade severely when deployed at real-world endpoints. Though training on data from endpoints leads to better performance, it still falls far short of expectations established in prior evaluations. We characterize various challenges ML approaches face in this domain and explore promising techniques targeting them, achieving moderate improvements. Ultimately, this task remains challenging. By exposing evaluation pitfalls and ML shortcomings, we call on the community to seek new solutions to these security-threatening problems.
X Acknowledgements
@@
This research was supported by the US Intelligence Community Postdoctoral Fellowship (Kaya), the US Department of Defense (Dumitraş), UK EPSRC Grant EP/X015971/1 (Pierazzi), US National Science Foundation Grants CNS-2154873 (Wagner) and CNS-2327427 (Botacin), and generous research awards from Google (Cavallaro and Wagner) and Amazon (Dumitraş and Wagner). We would like to thank VirusTotal for granting us academic access to their platform. Finally, we thank Omer Faruk Akgul and David Acs for their support in the earlier phases of our research and Erin Avllazagaj for providing us access to their program traces in the wild dataset. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the supporting organizations.
@@ Appendix A Trace Standardization
A-A Technical Details
Our EP traces contain high-level information the vendor can seamlessly collect and analyze at endpoint hosts, including file, process, and mutex creations, registry key creations and deletions, and process injections. Sandboxes also record lower-level information (such as memory dumps [86]) as they are less computationally constrained. In a standardized trace, first, we only keep the action types that all formats in our datasets share: file creation, registry key creation and deletion, process creation and injection, and mutex creation. Second, we clean up the strings (e.g., file names) in each trace by removing white spaces, capitalization, punctuation, non-ASCII characters, and so on. Third, we replace certain file and directory names to minimize differences caused by operating system versions or logging conventions. Fourth, we use regular expressions to replace MAC addresses, Windows security and resource identifiers, hashes (SHA-256, SHA-1, MD5), and epoch timestamps with special tokens (e.g. <macaddress>), to prevent introducing artifacts to our ML models. Fifth, we tokenize the entries in each trace (e.g., split a full file path into its components) and save a trace as a sequence of tokens, following [6]. We will release our sandbox dataset in this standardized format. Moreover, this format enables researchers to submit their detectors to be evaluated against our endpoint dataset (detailed in Section VIII) by providing them with an expected input format their detectors should accept. When we train a model, we only keep the top-10K tokens in terms of frequency in the training traces and replace the remaining tokens with special tokens, such as <rare_file_name>, again following [6, 7]. This makes the task more suitable for learning by eliminating uninformative tokens and reducing the feature dimensionality.
A-B The Performance Impact
We run the following experiment to understand whether our standardization hurts detection performance. We trained NGR models on unstandardized SB1 traces, which contain more action types (e.g., file reads, registry reads, and modules loaded) than standardized traces. We also kept the top-25K tokens (instead of the top-10K) from these traces for tokenization. The best model achieves 97.8% TPR on the SB1 test set (up from 95.0% when trained on standardized traces). However, its performance is 53.4% TPR on the SB2 test set (down from 60.7% when trained on standardized traces). This suggests that trace standardization moderately hurts the in-domain performance (on the test traces from the training sandbox) while improving generalization to other domains (to traces from other sandboxes). We hypothesize that standardization reduces overfitting to domain-specific features (discussed in Section VII-C), which is desirable in our study as we aim to maximize the performance of a sandbox-based model on endpoint traces. Note that, even after removing some features from sandbox traces, the classifier’s performance is still competitive with prior work (e.g., Neurlux [6]) that uses the full sandbox report. This is because different features (e.g., registry reads and registry creations) are already correlated and have redundancy, making each additional feature less effective.
@@ Appendix B Further Machine Learning Details
NGR (based on Maldy [5]). We convert each trace into a list of 2-grams. For example, the file path a/b/c.jpg is turned into three 2-grams: <a/b>, <b/c> and <c.jpg>. As this results in an intractable number of unique 2-grams (mostly very rare), we apply the hashing trick [87] that assigns a numerical value up to $2^{14}$ to each 2-gram. The final feature vector—$x$—for a trace is a $2^{14}$-dimensional vector and each dimension is set to the number of occurrences of the corresponding 2-gram in the trace. We use a ResNet-based architecture [88] to train on these features.
HYB (based on Neurlux [6]). We treat a trace as a natural language document and train an attention-convolution-hybrid sequence classification model. This approach eliminates the need for feature engineering (unlike the n-gram approach) and can extract useful features from long sequences thanks to the attention mechanism.
ATT (based on Nebula [7]). We treat a trace the same way as HYB (a sequence) and train a self-attention transformer-based architecture that is claimed to be robust to heterogeneous information (  e.g., different report format).
@@
Appendix C How to Do Model Selection in SB->EP
Following our observation in Section V-A, in this section, we assess whether finding a better model for SB->EP is possible by selecting the models using traces from an unseen sandbox (not used for training). We rank our models (100+ in each setting) based on their SB1, SB2, and EP performances (TPR) and compute Spearman’s correlation coefficients between these rankings. Table XI reveals that (except for one setting) (i) the rankings based on the training sandbox correlate poorly (sometimes negatively) with EP rankings, (ii) rankings based on an unseen sandbox correlate more strongly with EP rankings. For example, for ATT trained on SB1, the rankings based on SB1 and SB2 have $0.03$ and $0.45$ correlation with the rankings based on EP, respectively. These results support our claim that it yields better results when model selection in SB->EP is performed using traces from a different, unseen sandbox that is not used for training.
TABLE XI: Ranking correlations of sandbox-trained models according to their performances (TPR) on different test sets.
Appendix D Case Studies on Sandbox-Specific Artifacts
@@
Here, we dive deeper into the sandbox-specific artifacts we identified in Table IX.
@@
SogouExplorer is a Chinese web browser that exists only in malware traces (100% MalR.) in SB1; whereas it does not exist in any SB2 traces and very few EP traces. The samples that interact with it mainly belong to families such as Sivis and Memery, all tagged as file infectors that attach their code to other programs. Considering that a Chinese vendor developed SB1, we believe they pre-installed this browser on their sandboxes to generate an analysis environment representative of Chinese hosts. This, however, creates features specific to SB1 as samples interact with the programs in the environment. Although this artifact exists in a few endpoint traces from hosts in China, its prevalence is almost zero.
@@
PersonalBankPortal, according to our research, is a program distributed by a Chinese bank to its customers. The samples that inject into this program belong to families such as Tinba and Ramnit, all considered as banking trojans that specifically ex-filtrate banking data. We believe the vendor pre-installs this program to lure malware samples into exhibiting their behaviors. Although this practice is useful for analyzing a sample [56] (and for SB->SB), it causes artifacts that are rarely observed in the wild.
@@
Among the artifacts found in SB2, Spotify is a popular music streaming service, and Python is the interpreter for Python programming language. Both programs are targeted and injected by file infectors, similar to SogouExplorer in SB1. These programs are much less prevalent in endpoint traces than in SB2 traces. We believe the SB2 vendor, based in the US, pre-installs them to create an environment representative of the hosts in the US.
Appendix E The Impact of Sandbox Evasion
Recent works have measured that 40-80% of malware uses at least one evasive technique [17] to avoid analysis or terminate early if it is running in a sandbox [78, 21]. Evasion makes sandbox traces dissimilar to endpoint traces. Although sandbox evasion is well understood, its implications for endpoint detection have not been measured.
We explore the impact of sandbox evasion. We use a standard heuristic and treat a sample as evasive if the number of actions it performs is too low [28, 16, 21]. We first find the malware families common between our EP and SB1 test sets (30 total). We then compute the average trace length of each common family using the traces of the samples belonging to it. We count only registry actions, as they can be recorded unambiguously, unlike actions such as process injections, which might have vendor-specific definitions. We then find the length differences between SB1 and EP traces of each family. For example, Wannacry and Gandcrypt have differences of $+11$ and $-6$, respectively. We then split the families into two sets: the 21 families whose SB traces are longer (e.g., Wannacry) are deemed less likely to be evasive, and the 9 families whose SB traces are shorter (e.g., Gandcrypt) are deemed more likely. The median length differences for the non-evasive and evasive families are $+4.4$ and $-4.3$, respectively, with a few outliers on both sides, e.g., $+81$ for Vobfus and $-85$ for Bypassuac. Finally, we measure the TPR of our model in SB1->SB1 and SB1->EP on these two malware sets individually while keeping the benign samples the same.
TABLE XII: The performance (TPR) on malw. families more (Evasive) or less (Non-Eva) likely to be evading sandboxes.
The trace length of a sample monotonically increases as its execution continues. This means the discrepancies between the execution durations in SB1 and endpoint hosts might confound our measurements. Before presenting our results, we confirm this is unlikely to be the case: 82% of our EP traces are from executions that lasted less than 90 seconds. Although we do not know the exact configuration of SB1, allowing one to two minutes of execution is standard for most sandboxes in practice [28].
Based on the results in Table XII, we observe: (i) SB1->EP performance is higher (by 3%) on non-evasive families than on evasive families; (ii) SB1->SB1 performance is significantly higher (by 7%) on evasive families than on non-evasive families. This suggests that the classifier exploits short traces (evasiveness) as a feature, which is beneficial for SB->SB, though it does not generalize to SB->EP. However, the non-trivial SB1->EP performance on evasive families might hint that there are still features useful for endpoint detection in the sandbox traces of evasive samples. Although we are limited to observational data, a controlled study on evasive malware to disentangle these features is a promising direction.
These plots support our hypothesis: the model predicts a higher score for the shortest traces in both test sets, i.e., it has a length bias. This bias leads to accurate predictions in the SB set, where very short traces are much more likely to be malware (the malware ratio on the leftmost side of the upper plot is high). However, this bias leads to inaccurate predictions in the EP set, where very short are not more likely to be malware. On the EP traces shorter than 10 actions ($\sim$20of all EP traces), the model achieves 15% TPR (vs. 20% when all traces are kept). Ultimately, trace length is a spurious correlation learned from sandbox traces that fails to generalize to endpoint traces. Evidence suggests this is introduced by evasive malware that tends to produce short sandbox traces in the same execution time. Methods preventing the model from learning such correlations [72, 77] offer a promising next step.
Figure 5: Comparing the correlations between trace length and malware-ness prediction scores of the model. Malware ratio is the ground truth ratio of malware traces of a certain length among all traces of that length in a dataset.